solve the quadatric equations by factoring
x^2 - 9x - 4 = 6

Answer by stanbon(75887)

solve the quadatric equations by factoring
x^2 - 9x - 4 = 6
----------------------
x^2-9x-10 = 0
(x-10)(x+1)=0
x=10 or x=-1
===================
Cheers,
Stan H.